OpenStudy (anonymous):

What is an equation of the line that is parallel to the x-axis and that passes through the point (5,3)? Thanks x

OpenStudy (anonymous):

Parallel to x axis implies slope = 0, Using y = mx + b y = b is the answer: So y = 3 is the equation..

OpenStudy (campbell_st):

 the general form is y = mx + b m = slope, b = y-intercept. from the drawing the slope is 0 so y = 0x + b or y = b find the y-intercept